# The impact of Covid-19 vaccination in Aotearoa New Zealand: a modelling study

### 18 Jan 2024 v1

@@ Discussion
Vaccination may have prevented other adverse outcomes, such as symptomatic disease leading to attendance at primary care, lost productivity, and incidence of Long Covid [26, 27]. However, we only focused on hospitalisations and deaths and did not consider other outcomes in the model due to a lack of relevant data.
@@
Adverse events following immunisation with COVID-19 vaccines in Aotearoa New Zealand are monitored by the statutory regulator Medsafe. Detailed reports are available at [28]. Serious adverse events are extremely rare; however we have not attempted a formal risk-benefit analysis here.
@@
A global modelling study estimated that vaccines prevented approximately 20 million deaths up to December 2021 [12], which equates to around $2.5$ deaths prevented per 1000 primary series doses. Another analysis estimated that vaccines prevented 105,900 deaths in England up to August 2021 [29], equivalent to $1.4$ per 1000 primary series doses. By comparison our results equate to between $0.54$ and $1.28$ deaths prevented per 1000 primary series doses (of which there were $8.45$ million during the simulation period).
@@
There are several reasons why the number of deaths prevented per dose would be lower in the context of our study. In most countries, the initial vaccine rollout prioritised high-risk groups and took place during a period of high global community transmission of the Alpha, Beta and Delta variants. It is to be expected that there are diminishing returns from expanding the rollout to lower-risk groups, even though these are almost certainly still net beneficial. The Omicron variant is less pathogenic than earlier variants of concern [30] and effective treatments have improved over time. Finally, although vaccine effectiveness against severe disease caused by Omicron has remained high, effectiveness against transmission is significantly reduced. Globally, as well as in Aotearoa New Zealand, cumulative attack rates were far higher by June 2023 than in December 2021.
@@
Counterfactual modelling is a valuable tool to understand and learn from the potential consequences of alternative scenarios with the benefit of hindsight provided by subsequent data [31, 32, 33]. However, it has important limitations.
@@
In reality, in a scenario with low vaccine coverage, the occurrence of higher hospitalisation and death rates would have prompted different policy actions and behavioural responses [34], which would have altered the transmission dynamics compared to the counterfactual scenarios presented here. Therefore our results should not be interpreted as predictions of what would have happened if Aotearoa New Zealand had not achieved as high a vaccine coverage as it did. Nevertheless, in the absence of vaccination, non-pharmaceutical interventions would only have delayed rather than prevented infections. It is likely that the large majority of unvaccinated individuals would eventually have been infected, meaning a comparable number of hospitalisations and deaths as in the modelled counterfactual would have accrued, albeit over a longer time period. Our results are therefore a valid estimate of the reduction in health burden that is directly attributable to vaccines. We also note that the need for a more stringent policy or behavioural response would have incurred additional direct and indirect costs.
@@
The model for transmission dynamics does not include ethnicity as a variable. In estimating the number of Māori deaths that could have been prevented if vaccination rates had been equitable, we assumed that the per capita death rate for a subgroup with a given vaccination rate would be equal to the per capita death rate if that vaccination rate was applied to the whole population. This is a simplistic assumption that ignores the differential effect of vaccination on transmission between and within ethnicities. Our estimates apply only to the period from 1 January 2022 to 30 June 2023. They do not include vaccine-preventable health impacts that disproportionately affected Māori during the 2021 Delta outbreak [35, 13, 36].
@@
We found that estimates for the number of preventable Māori hospitalisations and deaths were larger when using StatsNZ population estimates than when using the HSU data for population denominators. It is recognised that the HSU data underestimates total Māori population size and has biases in the age and sex structure as it only includes those with recent engagement with health services [55]. Therefore the results that use the StatsNZ population estimates may be more accurate for Māori impacts. The significant gap between these point to the need for consistent collection of high-quality ethnicity data, accurate population size estimates, and explicit consideration of the uncertainty due to the differences in key data sources.
@@
Our results show that inequitable vaccine coverage is one contributing factor to the higher Covid-19 hospitalisation and mortality rates experienced by Māori. Other factors likely include increased exposure to infection due to high rates of employment in high-contact occupations, larger average household size and lower-quality housing [37, 38], high rates of comorbidities [9, 39, 40], poor access to healthcare [7], and a health system that has historically underserved Māori [8]. All these factors are worthy of attention and further study. Including the effects of ethnicity in dynamic epidemiological models is challenging due to the paucity of data and need for modelling assumptions about contact rates between and within ethnic groups. This a key area needing further research in Aotearoa New Zealand.
@@
Our results cannot be used to estimate outcomes if there had been widespread transmission of SARS-CoV-2 in Aotearoa New Zealand earlier in the pandemic when fewer effective treatments were available. Some pre-Omicron variants were more pathogenic [30, 41] and it is likely that health impacts would have been larger if uncontrolled transmission of these variants had occurred in an unvaccinated population.
@@
Our analysis is retrospective and cannot be used to predict the health benefits of future booster rollouts or changes to vaccine eligibility. These will depend on the future prevalence and clinical severity of SARS-CoV-2 variants, the marginal increase in protection offered by additional doses relative to existing levels of hybrid immunity, and how well matched the vaccines are to dominant circulating strains [42, 43].
@@
Aotearoa New Zealand’s Covid-19 pandemic response used a combination of border controls, physical distancing measures including periods of stay-at-home orders, financial support for those unable to work, testing and contact tracing, mass masking, vaccination, and antiviral treatments [1, 44, 45]. This response delivered one of the lowest pandemic mortality rates of any country in the world [46]. Given that non-pharmaceutical interventions could only delay rather than prevent widespread transmission, it is clear that vaccination was an essential component of this response that saved thousands of lives. Immunity provided by existing vaccine coverage and future vaccination will continue to be an important factor in reducing the Covid-19 health burden. This means the number of lives saved by Covid-19 vaccines in Aotearoa New Zealand will further increase over time. Ensuring equitable access to vaccination and vaccine coverage should be a priority for future public health or pandemic responses.

### 26 Jan 2024 v2

@@ 4 Discussion
Vaccination may have prevented other adverse outcomes, such as symptomatic disease leading to attendance at primary care, lost productivity, and incidence of Long Covid notarte2022impact ; tran2023efficacy . However, we only focused on hospitalisations and deaths and did not consider other outcomes in the model due to a lack of relevant data.
@@
Adverse events following immunisation with COVID-19 vaccines in Aotearoa New Zealand are monitored by the statutory regulator Medsafe. Detailed reports are available at medsafe_adverse . Serious adverse events are extremely rare; however we have not attempted a formal risk-benefit analysis here.
@@
A global modelling study estimated that vaccines prevented approximately 20 million deaths up to December 2021 watson2022global , which equates to around $2.5$ deaths prevented per 1000 primary series doses. Another analysis estimated that vaccines prevented 105,900 deaths in England up to August 2021 ukhsa2021vaccine , equivalent to $1.4$ per 1000 primary series doses. By comparison our results equate to between $0.54$ and $1.28$ deaths prevented per 1000 primary series doses (of which there were $8.45$ million during the simulation period).
@@
There are several reasons why the number of deaths prevented per dose would be lower in the context of our study. In most countries, the initial vaccine rollout prioritised high-risk groups and took place during a period of high global community transmission of the Alpha, Beta and Delta variants. It is to be expected that there are diminishing returns from expanding the rollout to lower-risk groups, even though these are almost certainly still net beneficial. The Omicron variant is less pathogenic than earlier variants of concern nyberg2022comparative and effective treatments have improved over time. Finally, although vaccine effectiveness against severe disease caused by Omicron has remained high, effectiveness against transmission is significantly reduced. Globally, as well as in Aotearoa New Zealand, cumulative attack rates were far higher by June 2023 than in December 2021.
@@
Counterfactual modelling is a valuable tool to understand and learn from the potential consequences of alternative scenarios with the benefit of hindsight provided by subsequent data mishra2021comparing ; binny2021early ; imai2023quantifying . However, it has important limitations.
@@
In reality, in a scenario with low vaccine coverage, the occurrence of higher hospitalisation and death rates would have prompted different policy actions and behavioural responses gimma2022changes , which would have altered the transmission dynamics compared to the counterfactual scenarios presented here. Therefore our results should not be interpreted as predictions of what would have happened if Aotearoa New Zealand had not achieved as high a vaccine coverage as it did. Nevertheless, in the absence of vaccination, non-pharmaceutical interventions would only have delayed rather than prevented infections. It is likely that the large majority of unvaccinated individuals would eventually have been infected, meaning a comparable number of hospitalisations and deaths as in the modelled counterfactual would have accrued, albeit over a longer time period. Our results are therefore a valid estimate of the reduction in health burden that is directly attributable to vaccines. We also note that the need for a more stringent policy or behavioural response would have incurred additional direct and indirect costs.
@@
The model for transmission dynamics does not include ethnicity as a variable. In estimating the number of Māori deaths that could have been prevented if vaccination rates had been equitable, we assumed that the per capita death rate for a subgroup with a given vaccination rate would be equal to the per capita death rate if that vaccination rate was applied to the whole population. This is a simplistic assumption that ignores the differential effect of vaccination on transmission between and within ethnicities. Our estimates apply only to the period from 1 January 2022 to 30 June 2023. They do not include vaccine-preventable health impacts that disproportionately affected Māori during the 2021 Delta outbreak tewhatuora_dashboard ; smith2022covid ; summers2023improvements .
@@
We found that estimates for the number of preventable Māori hospitalisations and deaths were larger when using StatsNZ population estimates than when using the HSU data for population denominators. It is recognised that the HSU data underestimates total Māori population size and has biases in the age and sex structure as it only includes those with recent engagement with health services statsnz_hsu . Therefore the results that use the StatsNZ population estimates may be more accurate for Māori impacts. The significant gap between these point to the need for consistent collection of high-quality ethnicity data, accurate population size estimates, and explicit consideration of the uncertainty due to the differences in key data sources.
@@
Our results show that inequitable vaccine coverage is one contributing factor to the higher Covid-19 hospitalisation and mortality rates experienced by Māori. Other factors likely include increased exposure to infection due to high rates of employment in high-contact occupations, larger average household size and lower-quality housing mcleod2020covid ; harvey2023summary , high rates of comorbidities steyn2020estimated ; yon2014cohort ; gurney2020inequity , poor access to healthcare whitehead2022structural , and a health system that has historically underserved Māori waitangi2021haumaru . All these factors are worthy of attention and further study. Including the effects of ethnicity in dynamic epidemiological models is challenging due to the paucity of data and need for modelling assumptions about contact rates between and within ethnic groups. This a key area needing further research in Aotearoa New Zealand.
@@
Our results cannot be used to estimate outcomes if there had been widespread transmission of SARS-CoV-2 in Aotearoa New Zealand earlier in the pandemic when fewer effective treatments were available. Some pre-Omicron variants were more pathogenic nyberg2022comparative ; davies2021increased and it is likely that health impacts would have been larger if uncontrolled transmission of these variants had occurred in an unvaccinated population. We did not attempt to model counterfactual scenarios based on alternative vaccination policies, such as absence of the vaccine pass system and employment-related vaccine mandates. This is because there is no quantitative or qualitative data to inform modelling assumptions about the effect these policies would have had on vaccine update or on contact rates of unvaccinated individuals.
@@
Our analysis is retrospective and cannot be used to predict the health benefits of future booster rollouts or changes to vaccine eligibility. These will depend on the future prevalence and clinical severity of SARS-CoV-2 variants, the marginal increase in protection offered by additional doses relative to existing levels of hybrid immunity, and how well matched the vaccines are to dominant circulating strains fabiani2023relative ; kirsebom2023effectiveness .
@@
Aotearoa New Zealand’s Covid-19 pandemic response used a combination of border controls, physical distancing measures including periods of stay-at-home orders, financial support for those unable to work, testing and contact tracing, mass masking, vaccination, and antiviral treatments baker2020successful ; cumming2022going ; binny2022real . This response delivered one of the lowest pandemic mortality rates of any country in the world cmi2023mortality . Given that non-pharmaceutical interventions could only delay rather than prevent widespread transmission, it is clear that vaccination was an essential component of this response that saved thousands of lives. Immunity provided by existing vaccine coverage and future vaccination will continue to be an important factor in reducing the Covid-19 health burden. This means the number of lives saved by Covid-19 vaccines in Aotearoa New Zealand will further increase over time. Ensuring equitable access to vaccination and vaccine coverage should be a priority for future public health or pandemic responses.